# Transferability of Adversarial Attacks in Video-based MLLMs: A Cross-modal Image-to-Video Approach

### v3

@@ 1 Introduction
Recent work has shown that video-based multimodal large language models (V-MLLMs) are vulnerable to adversarial video samples (li2024fmmattackflowbasedmultimodaladversarial), even though they have achieved remarkable performance on a wide range of video-text multimodal tasks (li2024videochatchatcentricvideounderstanding; jin2024chatuniviunifiedvisualrepresentation; lin2024videollavalearningunitedvisual; maaz2024videochatgptdetailedvideounderstanding; dai2023instructblipgeneralpurposevisionlanguagemodels; zhang2023videollamainstructiontunedaudiovisuallanguage). Existing work primarily focuses on white-box attacks, where information about the target model is accessible. However, the transferability of adversarial video samples across V-MLLMs remains unexplored, which is a more common and practical setting in real-world scenarios. It is still uncertain whether the adversarial videos generated on the source model can effectively attack other target models, posing significant security risks to the deployment of V-MLLMs in real-world applications.
In this paper, we pioneer an investigation into the transferability of adversarial video samples across V-MLLMs. Through detailed analysis in Sec. 3.2, we think previous methods have these shortcomings: (1) lacking generalization in perturbing video features, (2) focusing only on sparse key-frames, and (3) failing to integrate multimodal information. FMM attack (li2024fmmattackflowbasedmultimodaladversarial) is the first proposed white-box attack method targeting V-MLLMs. It utilizes flow-based temporal mask to select key-frames and applies perturbations to these frames. FMM attack performs well in the white-box setting but has limited transferability in the black-box setting. FMM attack heavily relies on the video features, which causes the generated perturbations to overfit to the video features extracted by the surrogate model, thereby reducing their generalizability. Additionally, since FMM attack applies perturbations only to key-frames, it cannot ensure that all frames sampled by the target model are perturbed. Taking low-level image features into account can help with improving transferability of adversarial samples. Previous image-to-video cross-modal attacks (wei2021crossmodaltransferableadversarialattacks; wang2023global; kim2023breakingtemporalconsistencygenerating) demonstrate the possibility of using image models as surrogates to attack video models in the black-box setting. However, these traditional attack methods typically focus on the video classification tasks with vision-only models, failing to integrate multimodal information.
@@
We conduct various experiments on three well-established datasets, MSVD-QA (xu2017video), MSRVTT-QA (xu2017video), and ActivityNet-200 (Heilbron_Escorcia_Ghanem_Niebles_2015) to evaluate the performance of our proposed I2V-MLLM attack in multiple video-text multimodal tasks. The experimental results demonstrate that our method can generate adversarial videos with strong transferability across different V-MLLMs (Chat-Univi (jin2024chatuniviunifiedvisualrepresentation), LLava-Next-Video (zhang2024videoinstructiontuningsynthetic), VideoChat, Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage)), and achieve competitive performance with white-box attacks against V-MLLMs. Our main contributions are summarized as follows:
@@ 2.1 Multimodal large language models
MLLMs typically consist of a vision model, a pretrained LLM, and a projector that translates visual information into textual representations that the LLM can process. Currently, MLLMs can be categorized into image-based and video-based types. I-MLLMs (dai2023instructblipgeneralpurposevisionlanguagemodels; liu2023visualinstructiontuning; zhu2023minigpt4enhancingvisionlanguageunderstanding; alayrac2022flamingovisuallanguagemodel; awadalla2023openflamingoopensourceframeworktraining; hu2023blivasimplemultimodalllm; bai2023qwenvlversatilevisionlanguagemodel) are designed to handle image-text inputs. They excel in tasks such as visual question answering, image captioning, and more. V-MLLMs extend the capabilities of I-MLLMs by incorporating temporal modules that allow them to understand and process video input. This enables them to perform tasks like video question answering (VideoQA), spatiotemporal localization, and video captioning. For example, Chat-UniVi (jin2024chatuniviunifiedvisualrepresentation) extracts specific frames from videos and utilizes DPC-KNN (Du_Ding_Jia_2016) to group these frames into distinct events, Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage) employs sequential encoding to capture temporal relationships among video frames, VideoChatGPT (maaz2024videochatgptdetailedvideounderstanding) applies temporal pooling to video features to extract temporal information. These methods equip the models with the capability to capture and interpret temporal dynamics, thus enabling a more comprehensive understanding of video content.
@@ 2.2 Adversarial attacks on MLLMs
Despite the impressive performance, MLLMs are highly susceptible to adversarial attacks (zhao2023evaluatingadversarialrobustnesslarge; luo2024imageworth1000lies; cui2023robustnesslargemultimodalmodels; tu2023many; zhang2024avibench; bailey2023image; lu2023set). For I-MLLMs, several studies have assessed their vulnerabilities to adversarial attacks. (fu2023misusingtoolslargelanguage) introduces Trojan-like images that force the target models to invoke malicious external tools or APIs specified by the attacker. (dong2023robustgooglesbardadversarial) utilizes open-source MLLMs to generate transferable adversarial examples capable of attacking closed-source commercial models like Bard (google2023gemini), Bing Chat (microsoft2024bingchat), and GPT-4V (openai2024gpt4v), thereby showing high transferability of adversarial examples across MLLMs. (gu2025improving) introduces DynVLA Attack, which applies dynamic perturbations to the vision-language connector to improve the generalization of adversarial attacks across different alignment strategies. While extensive studies have explored adversarial attacks on I-MLLMs, there has been little exploration in the domain of V-MLLMs. (li2024fmmattackflowbasedmultimodaladversarial) proposes a flow-based adversarial attack strategy for white-box attacks on V-MLLMs. However, in real-world scenarios, the internal architectures and parameters of V-MLLMs are usually inaccessible to users. To address this, we focus on exploring methods for conducting adversarial attacks on V-MLLMs in a black-box setting.
@@ 2.3 Adversarial attacks on video models
Current video models have diverse applications, including autonomous vehicles, video verification, security, and other fields. However, these models remain vulnerable to adversarial attacks (li2018adversarial; xie2022universal; jiang2019black; wei2020heuristic; cao2024logostylefoolvitiatingvideorecognition). For example, Universal 3D perturbations (U3D) (xie2022universal) deceive video classifiers by generating a universal perturbation for all input videos, while StyleFool (cao2023stylefool) introduces an unrestricted perturbation to attack video classification systems through style transfer. Recent studies also explore cross-modal attack methods from image models to video models (wei2021crossmodaltransferableadversarialattacks; kim2023breakingtemporalconsistencygenerating; wang2023global), yielding promising results. However, these attacks primarily target video classification tasks, which do not account for interactions between visual and textual modalities. In contrast, V-MLLMs integrate both visual and textual information, rendering these methods unsuitable for such models. To tackle this issue, our method introduces multimodal interactions in adversarial video crafting, aligning seamlessly with how V-MLLMs operate.
@@ 3.1 Preliminary
Given a video sample $V\in\mathcal{V}$ with $M$ associated QA pairs $\{(q_{m},a_{m})\}_{m=1}^{M}$, where $q_{m}$ is the $m$-th question and $a_{m}$ is the corresponding answer. Let $F$ denote the I-MLLM (e.g., BLIP-2 (li2023blip2bootstrappinglanguageimagepretraining), MiniGPT-4 (zhu2023minigpt4enhancingvisionlanguageunderstanding)) and $G$ denote the V-MLLM (e.g., Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage), Chat-UniVi. We use $G(V,q)$ to denote the answer generated by the V-MLLM for the given video $V$ and question $q$. The goal of our proposed attack is to generate an adversarial example $V_{adv}=V+\delta^{\prime}$ using $F$, which can cause $G$ to produce an answer $G(V_{adv},q_{i})$ that differs significantly from the correct answer $a_{i}$, without accessing the parameters or structure of $G$, where $\delta^{\prime}$ denotes the adversarial perturbations specifically tailored for $V$. To ensure that the adversarial perturbation $\delta^{\prime}$ is imperceptible, we restrict it by $\delta^{\prime}||_{\infty}\leq\epsilon$, where $||\cdot||_{\infty}$ denotes the $L_{\infty}$ norm, and $\epsilon$ is a constant for the norm constraint. We utilize the evaluation model $E$ (i.e., GPT-4o-mini (openai2023gpt4omini)) to assess whether the generated answer aligns with the reference answer. We aim to find imperceptible adversarial perturbations that minimize the number of correct responses, formulated as follows:
@@ 3.2 Motivation
Lacking generalization in perturbing video features. Adversarial perturbations generated based on certain V-MLLM can overfit to specific video module, limiting their generalization to other V-MLLMs. To improve transferability, we focus on lower-level image features. The I2V attack (wei2021crossmodaltransferableadversarialattacks), which perturbs each video frame to disrupt image features, demonstrates improved transferability when using image models as surrogates to craft adversarial video samples, as shown in rows 3, 4, and 5 of Tab. 1.
@@
Table 1: Attack success rates (ASR, %) on the MSVD-QA validation set for Zero-Shot VideoQA tasks. FMM and I2V denote attack methods from (li2024fmmattackflowbasedmultimodaladversarial) and (wei2021crossmodaltransferableadversarialattacks), respectively. Vanilla attack applies full perturbations on all key-frames sampled by V-MLLMs. Prop. denotes perturbation propagation. * indicates white-box attacks. AASR represents the average ASR across all target models for each surrogate model. A higher AASR indicates better adversarial transferability.
@@ 3.3 I2V-MLLM Attack
To further disrupt V-MLLMs’ capacity for video-text multimodal tasks, I2V-MLLM interferes with the intermediate feature of the projector (e.g. Q-Former (li2023blip2bootstrappinglanguageimagepretraining)), which plays an essential role in aligning visual and textual representations. We feed the projector with the original frame-level embeddings $F_{V}(X)$, the adversarial frame-level embeddings $F_{V}(X_{adv})$ from the vision model, and the caption set $T=\{t_{1},t_{2},\ldots,t_{M}\}$. After multimodal alignment, they are transformed into the original visual features $F_{P}^{v}(X)\in\mathbb{R}^{N_{1}\times D_{2}}$, the adversarial visual features $F_{P}^{v}(X_{adv})\in\mathbb{R}^{N_{1}\times D_{2}}$, and the textual features $F_{P}^{t}(T)\in\mathbb{R}^{N_{2}\times D_{2}}$. Here, $N_{1}$ and $N_{2}$ represent the number of visual features and the textual features, respectively. And $D_{2}$ denotes the dimension of these features. The captions are complete sentences generated based on the question $q$ and the answer $a$ using GPT-4o-mini (openai2023gpt4omini). For example, given the question $q$: ‘What is the man doing?’ and the answer $a$: ‘eat’, the corresponding caption $t$ would be: ‘The man is eating.’ To perturb the image features aligned with the text, I2V-MLLM optimizes the adversarial perturbations by minimizing the cosine similarity between the original and the adversarial visual features:
@@ 4.1 Experimental setting
Datasets and models. Referring to the quantitative benchmarking framework proposed in (maaz2024videochatgptdetailedvideounderstanding), we evaluate our I2V-MLLM attack on Zero-Shot VideoQA tasks using the validation set of MSVD-QA and MSRVTT-QA. We perform the proposed method on three I-MLLMs: BLIP-2, InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels), and MiniGPT-4. Our method is evaluated on four different V-MLLMs: Chat-UniVi, LLaVA-NeXT-Video, VideoChat, and Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage), each with a Vicuna-7B (vicuna2023) as the LLM.
Attack setting. In I2V-MLLM, we employ the projected gradient descent (PGD) (madry2019deeplearningmodelsresistant) with a perturbation bound of $\epsilon=16$, an iteration number of $I=50$, and a step size of $\alpha=1$ for the attack process. The parameters $\lambda_{1}$ and $\lambda_{2}$ are both set to $1$, and the key-frame ratio $\beta$ is set to $30\%$. I2V attack, utilizing CLIP-L/14 (radford2021learningtransferablevisualmodels) as the surrogate model, applies tailored perturbations to each frame of the video. For a fair comparison, the PGD parameters ($\epsilon=16$, $I=50$ and $\alpha=1$) in FMM, Vanilla, and I2V attacks maintain the same for our method. Additionally, in the FMM setup, the key-frame ratio $\beta$ is also set to $30\%$. All the experiments are conducted on an NVIDIA-A6000 GPU.
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on Zero-Shot VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini, which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods (maaz2024videochatgptdetailedvideounderstanding) to assess Accuracy (Acc.) and GPT-Score. Specifically, accuracy (Acc.) refers to the model’s prediction accuracy, while the GPT score (Score) assesses the quality of the model’s predictions, assigning a relative score on a scale from 0 to 5. GPT-4o-mini is used for evaluation due to its strong text understanding and cost efficiency. For detailed explanations of the metrics, see Appendix B.1.

### v4

@@ 1 Introduction
Recent work has shown that video-based multimodal large language models (V-MLLMs) are vulnerable to adversarial video samples (Li et al. 2024a), even though they have achieved remarkable performance on a wide range of video-text multimodal tasks (Li et al. 2024b; Jin et al. 2024; Lin et al. 2024; Maaz et al. 2024; Dai et al. 2023; Zhang et al. 2023). Existing work primarily focuses on white-box attacks, where information about the target model is accessible. However, the transferability of adversarial video samples across V-MLLMs remains unexplored, which is a more common and practical setting in real-world scenarios. It is still uncertain whether the adversarial videos generated on the source model can effectively attack other target models, posing significant security risks to the deployment of V-MLLMs in real-world applications.
In this paper, we pioneer an investigation into the transferability of adversarial video samples across V-MLLMs. Through detailed analysis in Sec. 3.2, we think previous methods have these shortcomings: (1) lacking generalization in perturbing video features, (2) focusing only on sparse key-frames, and (3) failing to integrate multimodal information. FMM attack (Li et al. 2024a) is the first proposed white-box attack method targeting V-MLLMs. It utilizes flow-based temporal mask to select key-frames and applies perturbations to these frames. FMM attack performs well in the white-box setting but has limited transferability in the black-box setting. FMM attack heavily relies on the video features, which causes the generated perturbations to overfit to the video features extracted by the surrogate model, thereby reducing their generalizability. Additionally, since FMM attack applies perturbations only to key-frames, it cannot ensure that all frames sampled by the target model are perturbed. Taking low-level image features into account can help with improving transferability of adversarial samples. Previous image-to-video cross-modal attacks (Wei et al. 2021; Wang et al. 2023; Kim et al. 2023) demonstrate the possibility of using image models as surrogates to attack video models in the black-box setting. However, these traditional attack methods typically focus on the video classification tasks with vision-only models, failing to integrate multimodal information.
@@
We conduct various experiments on three well-established datasets, MSVD-QA (Xu et al. 2017), MSRVTT-QA (Xu et al. 2017), and ActivityNet-200 (Heilbron et al. 2015) to evaluate the performance of our proposed I2V-MLLM attack in multiple video-text multimodal tasks. The experimental results demonstrate that our method can generate adversarial videos with strong transferability across different V-MLLMs (Chat-Univi (Jin et al. 2024), LLava-Next-Video (Zhang et al. 2024b), VideoChat, Video-LLaMA (Zhang et al. 2023)), and achieve competitive performance with white-box attacks against V-MLLMs. Our main contributions are summarized as follows:
@@ 2.1 Multimodal large language models
MLLMs typically consist of a vision model, a pretrained LLM, and a projector that translates visual information into textual representations that the LLM can process. Currently, MLLMs can be categorized into image-based and video-based types. I-MLLMs (Dai et al. 2023; Liu et al. 2023; Zhu et al. 2023; Alayrac et al. 2022; Awadalla et al. 2023; Hu et al. 2023; Bai et al. 2023) are designed to handle image-text inputs. They excel in tasks such as visual question answering, image captioning, and more. V-MLLMs extend the capabilities of I-MLLMs by incorporating temporal modules that allow them to understand and process video input. This enables them to perform tasks like video question answering (VideoQA), spatiotemporal localization, and video captioning. For example, Chat-UniVi (Jin et al. 2024) extracts specific frames from videos and utilizes DPC-KNN (Du et al. 2016) to group these frames into distinct events, Video-LLaMA (Zhang et al. 2023) employs sequential encoding to capture temporal relationships among video frames, VideoChatGPT (Maaz et al. 2024) applies temporal pooling to video features to extract temporal information. These methods equip the models with the capability to capture and interpret temporal dynamics, thus enabling a more comprehensive understanding of video content.
@@ 2.2 Adversarial attacks on MLLMs
Despite the impressive performance, MLLMs are highly susceptible to adversarial attacks (Zhao et al. 2023; Luo et al. 2024; Cui et al. 2023; Tu et al. 2023; Zhang et al. 2024a; Bailey et al. 2023; Lu et al. 2023). For I-MLLMs, several studies have assessed their vulnerabilities to adversarial attacks. (Fu et al. 2023) introduces Trojan-like images that force the target models to invoke malicious external tools or APIs specified by the attacker. (Dong et al. 2023) utilizes open-source MLLMs to generate transferable adversarial examples capable of attacking closed-source commercial models like Bard (Google 2024), Bing Chat (Microsoft 2024), and GPT-4V (OpenAI 2024b), thereby showing high transferability of adversarial examples across MLLMs. (Gu et al. 2025) introduces DynVLA Attack, which applies dynamic perturbations to the vision-language connector to improve the generalization of adversarial attacks across different alignment strategies. While extensive studies have explored adversarial attacks on I-MLLMs, there has been little exploration in the domain of V-MLLMs. (Li et al. 2024a) proposes a flow-based adversarial attack strategy for white-box attacks on V-MLLMs. However, in real-world scenarios, the internal architectures and parameters of V-MLLMs are usually inaccessible to users. To address this, we focus on exploring methods for conducting adversarial attacks on V-MLLMs in a black-box setting.
@@ 2.3 Adversarial attacks on video models
Current video models have diverse applications, including autonomous vehicles, video verification, security, and other fields. However, these models remain vulnerable to adversarial attacks (Li et al. 2018; Xie et al. 2022; Jiang et al. 2019; Wei et al. 2020; Cao et al. 2024). For example, Universal 3D perturbations (U3D) (Xie et al. 2022) deceive video classifiers by generating a universal perturbation for all input videos, while StyleFool (Cao et al. 2023) introduces an unrestricted perturbation to attack video classification systems through style transfer. Recent studies also explore cross-modal attack methods from image models to video models (Wei et al. 2021; Kim et al. 2023; Wang et al. 2023), yielding promising results. However, these attacks primarily target video classification tasks, which do not account for interactions between visual and textual modalities. In contrast, V-MLLMs integrate both visual and textual information, rendering these methods unsuitable for such models. To tackle this issue, our method introduces multimodal interactions in adversarial video crafting, aligning seamlessly with how V-MLLMs operate.
@@ 3.1 Preliminary
Given a video sample $V\in\mathcal{V}$ with $M$ associated QA pairs $\{(q_{m},a_{m})\}_{m=1}^{M}$, where $q_{m}$ is the $m$-th question and $a_{m}$ is the corresponding answer. Let $F$ denote the I-MLLM (e.g., BLIP-2 (Li et al. 2023), MiniGPT-4 (Zhu et al. 2023)) and $G$ denote the V-MLLM (e.g., Video-LLaMA (Zhang et al. 2023), Chat-UniVi. We use $G(V,q)$ to denote the answer generated by the V-MLLM for the given video $V$ and question $q$. The goal of our proposed attack is to generate an adversarial example $V_{adv}=V+\delta^{\prime}$ using $F$, which can cause $G$ to produce an answer $G(V_{adv},q_{i})$ that differs significantly from the correct answer $a_{i}$, without accessing the parameters or structure of $G$, where $\delta^{\prime}$ denotes the adversarial perturbations specifically tailored for $V$. To ensure that the adversarial perturbation $\delta^{\prime}$ is imperceptible, we restrict it by $\delta^{\prime}||_{\infty}\leq\epsilon$, where $||\cdot||_{\infty}$ denotes the $L_{\infty}$ norm, and $\epsilon$ is a constant for the norm constraint. We utilize the evaluation model $E$ (i.e., GPT-4o-mini (OpenAI 2024a)) to assess whether the generated answer aligns with the reference answer. We aim to find imperceptible adversarial perturbations that minimize the number of correct responses, formulated as follows:
@@ 3.2 Motivation
Lacking generalization in perturbing video features. Adversarial perturbations generated based on certain V-MLLM can overfit to specific video module, limiting their generalization to other V-MLLMs. To improve transferability, we focus on lower-level image features. The I2V attack (Wei et al. 2021), which perturbs each video frame to disrupt image features, demonstrates improved transferability when using image models as surrogates to craft adversarial video samples, as shown in rows 3, 4, and 5 of Tab. 1.
@@
Table 1: Attack success rates (ASR, %) on the MSVD-QA validation set for Zero-Shot VideoQA tasks. FMM and I2V denote attack methods from (Li et al. 2024a) and (Wei et al. 2021), respectively. Vanilla attack applies full perturbations on all key-frames sampled by V-MLLMs. Prop. denotes perturbation propagation. * indicates white-box attacks. AASR represents the average ASR across all target models for each surrogate model. A higher AASR indicates better adversarial transferability.
@@ 3.3 I2V-MLLM Attack
To further disrupt V-MLLMs’ capacity for video-text multimodal tasks, I2V-MLLM interferes with the intermediate feature of the projector (e.g. Q-Former (Li et al. 2023)), which plays an essential role in aligning visual and textual representations. We feed the projector with the original frame-level embeddings $F_{V}(X)$, the adversarial frame-level embeddings $F_{V}(X_{adv})$ from the vision model, and the caption set $T=\{t_{1},t_{2},\ldots,t_{M}\}$. After multimodal alignment, they are transformed into the original visual features $F_{P}^{v}(X)\in\mathbb{R}^{N_{1}\times D_{2}}$, the adversarial visual features $F_{P}^{v}(X_{adv})\in\mathbb{R}^{N_{1}\times D_{2}}$, and the textual features $F_{P}^{t}(T)\in\mathbb{R}^{N_{2}\times D_{2}}$. Here, $N_{1}$ and $N_{2}$ represent the number of visual features and the textual features, respectively. And $D_{2}$ denotes the dimension of these features. The captions are complete sentences generated based on the question $q$ and the answer $a$ using GPT-4o-mini (OpenAI 2024a). For example, given the question $q$: ‘What is the man doing?’ and the answer $a$: ‘eat’, the corresponding caption $t$ would be: ‘The man is eating.’ To perturb the image features aligned with the text, I2V-MLLM optimizes the adversarial perturbations by minimizing the cosine similarity between the original and the adversarial visual features:
@@ 4.1 Experimental setting
Datasets and models. Referring to the quantitative benchmarking framework proposed in (Maaz et al. 2024), we evaluate our I2V-MLLM attack on Zero-Shot VideoQA tasks using the validation set of MSVD-QA and MSRVTT-QA. We perform the proposed method on three I-MLLMs: BLIP-2, InstructBLIP (Dai et al. 2023), and MiniGPT-4. Our method is evaluated on four different V-MLLMs: Chat-UniVi, LLaVA-NeXT-Video, VideoChat, and Video-LLaMA (Zhang et al. 2023), each with a Vicuna-7B (Chiang et al. 2023) as the LLM.
Attack setting. In I2V-MLLM, we employ the projected gradient descent (PGD) (Madry et al. 2019) with a perturbation bound of $\epsilon=16$, an iteration number of $I=50$, and a step size of $\alpha=1$ for the attack process. The parameters $\lambda_{1}$ and $\lambda_{2}$ are both set to $1$, and the key-frame ratio $\beta$ is set to $30\%$. I2V attack, utilizing CLIP-L/14 (Radford et al. 2021) as the surrogate model, applies tailored perturbations to each frame of the video. For a fair comparison, the PGD parameters ($\epsilon=16$, $I=50$ and $\alpha=1$) in FMM, Vanilla, and I2V attacks maintain the same for our method. Additionally, in the FMM setup, the key-frame ratio $\beta$ is also set to $30\%$. All the experiments are conducted on an NVIDIA-A6000 GPU.
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on Zero-Shot VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini, which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods (Maaz et al. 2024) to assess Accuracy (Acc.) and GPT-Score. Specifically, accuracy (Acc.) refers to the model’s prediction accuracy, while the GPT score (Score) assesses the quality of the model’s predictions, assigning a relative score on a scale from 0 to 5. GPT-4o-mini is used for evaluation due to its strong text understanding and cost efficiency. For detailed explanations of the metrics, see Appendix B.1.